OpenStudy (anonymous):

n --- =300 12 n = a0 plez help me:(

OpenStudy (raden):

cross multiply to get n

OpenStudy (anonymous):

wait can u remind me how to do that becuzz i forgpot the order

OpenStudy (raden):

n/12 = 300 n = 12 * 300 = ...

OpenStudy (texaschic101):

n/12 = 300/1 (cross multiply) (n)(1) = (300)(12) n = 300 x 12 (you can finish this)

OpenStudy (texaschic101):

Oh my....I just realized this question is 2 month old. Why is it still open ?